OpenStudy (anonymous):

pls. help me simplifying this: y-1/(2y-1)y-1

OpenStudy (anonymous):

\[\frac{y-1}{(2y-1)(y-1)}\]?

OpenStudy (anonymous):

the y-1 in numerator and denominator will cancel out

OpenStudy (anonymous):

if so get \[\frac{1}{2y-1}\] but i am not sure that is what you meant.

OpenStudy (dumbcow):

\[\frac{y-1}{(y-1)(2y-1)}\] the y-1 cancels on top and bottom because they are like factors (y-1)/(y-1) = 1 \[\frac{y-1}{(y-1)(2y-1)}= \frac{1}{2y-1} \]

OpenStudy (anonymous):

\[y-1/(2y-1)y-1 \]

OpenStudy (anonymous):

thats the equation